# Analogical Structure, Minimal Contextual Cues and Contrastive Distractors: Input Design for Sample-Efficient Linguistic Rule Induction

### v1

@@ Abstract
Large language models achieve strong performance through training on vast datasets. Can analogical paradigm organization enable lightweight models to match this performance with minimal data? We develop a computational approach implementing three cognitive-inspired principles: analogical structure, contrastive learning, and minimal contextual cues. We test this approach with structured completion tasks where models identify correct sentence completions from analogical patterns with contrastive alternatives. Training lightweight models (BERT+CNN, $0.5M$ parameters) on only one hundred structured examples of English causative/inchoative alternations achieves $F1=0.95$, outperforming zero-shot GPT-o3 ($F1=0.87$). Ablation studies confirm that analogical organization and contrastive structure improve performance, consistently surpassing randomly shuffled baselines across architectures. Cross-phenomenon validation using unspecified object alternations replicates these efficiency gains, confirming approach robustness. Our results show that analogical paradigm organization enables competitive linguistic rule learning with orders of magnitude less data than conventional approaches require.
@@ 1 Introduction
Analogical reasoning, recognizing structural relationships across different contexts, enables efficient learning by abstracting patterns from minimal examples and transferring knowledge to novel situations (lake2017building). However, this principle has faced persistent scalability challenges across computational approaches: early symbolic systems required extensive knowledge engineering (forbus1995mac), neural word embeddings showed inconsistent performance (mikolov2013efficient), and even transformer-based models like BERT show mixed capabilities (ushio-etal-2021-bert; thrush2020investigating). While recent systems achieve strong performance (yasunaga2024large; jiayang-etal-2023-storyanalogy; wijesiriwardene-etal-2023-analogical), they still require substantial computational resources or show inconsistent results across complexity levels.
Rather than scaling analogical reasoning systems, we operationalise analogical principles through strategic input organization. Research shows that processing constraints can optimize attention allocation (christiansen2016now), while contrastive learning frameworks show how systematic positive-negative comparisons improve discriminative learning (chen2020simple; he2020momentum). This suggests that structural organization, not architectural scaling, may unlock analogical efficiency.
@@
In this paper, we develop a computational approach that organizes input data into analogical paradigms through three cognitive-inspired strategies: (i) Analogical Structure, systematic paradigmatic relationships supporting pattern recognition; (ii) Contrastive Learning, systematically designed distractors enabling discriminative learning; and (iii) Minimal Contextual Cues, subtle annotations providing essential semantic information without explicit labeling.
We evaluate this setup using English verb alternations (Levin93) through structured completion tasks based on Blackbird Language Matrices (merlo-2023-blackbird). These tasks require identifying correct sentence completions from systematically organized contexts with contrastive alternatives. Figure 1 illustrates our approach: contexts encode analogical mappings across parallel paradigms (e.g., Man:Dice :: Exploring:Mat), requiring models to abstract structural and semantic role relationships rather than memorize surface forms.
Results demonstrate sample efficiency: lightweight models (BERT+CNN, $\sim 0.5M$ parameters) trained on only $100$ structured examples achieve $F1=0.95$, outperforming zero-shot GPT-o3 ($F1=0.87$). Cross-phenomenon validation confirms robustness beyond causative constructions.
Scope: We operationalise computational algorithms inspired by cognitive principles without claiming to model human cognition directly.
Our contributions are: (i) a computational approach that implements analogical paradigm organization for sample-efficient linguistic rule learning; (ii) empirical evidence that this approach enables competitive performance with minimal training examples; (iii) systematic demonstration that lightweight models can match zero-shot LLM performance on linguistic rule induction tasks.  [^1] Code, data, and replication steps will be made publicly available upon publication.
@@ 2 Data and Task Design
We operationalize three cognitive-inspired principles through systematic paradigm organization using Blackbird Language Matrices (merlo-2023-blackbird), structured completion tasks that test systematic linguistic rule learning. This section describes our task framework, linguistic domain, and distractor design methodology.
@@ 2.1 Task Framework
Our completion tasks organize linguistic data into analogical paradigm structures. Each instance contains a structured Context comprising two parallel paradigms encoding analogical relationships (Pattern A $\leftrightarrow$ Pattern B), supplemented with minimal contextual cues that provide semantic scaffolding without explicit labeling. It also contains a systematic Answer Set: One correct completion that maintains established patterns, and six systematically designed distractors that violate specific structural, semantic, or paradigmatic constraints.
Success requires integrating our three cognitive principles. Models must recognize analogical mappings across paradigms (analogical structure), discriminate between valid and invalid transformations (contrastive learning), and use semantic scaffolding from minimal soft annotations (minimal contextual cues).
@@ 2.2 Linguistic Test Domain: English Verb Alternations
We validate the generalisability of the approach with bake-class verbs, a class of verbs exhibiting the unspecified object alternations (The chef baked a cake vs. The chef baked). In this class, the subject of both the transitive and the intransitive is an Agent. While the two classes share syntactic structure, they instantiate semantically distinct processes, testing whether our organizational principles generalize beyond specific semantic classes.
2.3 Analogical Paradigm Organization
Figure 1 illustrates our systematic organization approach. Each context follows a $2\times 4$ paradigmatic structure.
@@ 2.3 Analogical Paradigm Organization
We provide semantic scaffolding through subtle annotations within each paradigm. For example, action descriptors indicate agentivity (The player did it) while state descriptions suggest thematic roles (The ball was on the floor). These cues clarify semantic relationships without explicit grammatical labeling.
@@
Each distractor violates exactly one constraint dimension while preserving others, enabling precise evaluation of rule component acquisition. Role errors (RR) test semantic understanding, structural errors (SC-RR, SCRS) test syntactic knowledge, and paradigm errors (PC-RR, PSC-RR, PSC-RS) test analogical consistency across contexts.
@@
2.4 Data Generation and Availability
Our systematic generation process uses template-based sentence creation with controlled lexical variation and seed alternation pairs, following established verb classifications (Levin93). [^2] Verb class and seed sentences are detailed in Appendix B
@@ 3 Experiments
We systematically evaluate our analogical paradigm organization approach through controlled experiments testing three core claims: (i) our organizational principles enable sample efficiency compared to unstructured baselines, (ii) individual components contribute systematically to performance, and (iii) lightweight models trained with structured data can match zero-shot reasoning-capable LLM performance on linguistic rule learning. The performance on BLM puzzle completion is used as a proxy for a model’s level of successful linguistic rule learning.
@@ 3.1 Experimental Design
We test five systematic conditions that isolate organizational components while controlling information content.
Base is the complete paradigmatic organization with analogical structure and contextual cues. Shuffled has identical content in random order, (tests organizational vs. content effects). NoAnalogy removes first paradigm (tests analogical structure contribution). NoSoftCue removes contextual annotations (tests scaffolding contribution). Transposed preserves analogical patterns but changes spatial arrangement ( $C_{\text{Trans}}=C_{\text{Base}}^{\top}$).  [^3] See the formal definitions in Appendix A.1.
@@
Type I uses same verbs across paradigms ($3000$ examples, 80:10:10 splits, direct analogical mapping). Type II uses different verbs across paradigms ($15000$ examples, same splits, tests abstraction). Cross-phenomenon uses unspecified object alternations with bake-class verbs (validates generalisability).
@@
Training sizes range from $10$ to $2700$ examples for lightweight models; and $300$-example test sets (same as for Type I test data, and $300$ random examples from Type II test set) for LLMs. This balances computational feasibility with statistical reliability for cross-model comparisons.
@@ 3.2 Models and Training
We use several models of representation. We embed each sentence using BERT444bert-base-multilingual-cased(devlin-etal-2019-bert) as primary model, RoBERTa555xlm-roberta-base (liu-etal-2019-roberta) and ELECTRA666electra-base-discriminator (clark-etal-2020-electra) for encoder comparison. We reason that using a weaker base clarifies contributions attributable to data organization rather than encoder capabilities.
As for architecture types, we use previously tested parameter-efficient neural models (CNN and FFNN in nastase-merlo-2023-grammatical for comparison).(1) CNN captures localized patterns and positional relationships; (2) FFNN processes concatenated embeddings through fully-connected layers, integrating distributed information across the entire context. Both architectures maintain minimal parameter counts ($\sim 0.5M$ parameters) compared to the encoder ($\sim 110M$), isolating the influence of data organization from architectural capacity. Both models output an answer embedding given the concatenated embeddings of the sentences from the context, evaluated against answer candidates using a max-margin objective with cosine similarity. Training: $120$ epochs, learning rate $0.001$, Adam Optimizer, batch size $100$, early stopping (patience $10$). All experiments use random seed $42$ with results averaged over $3$ independent runs. More model and training details are provided in the Appendix.
@@
We evaluate eight advanced models including reasoning-capable systems. Reasoning models: deepseek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability), gpt-o3, gpt-o3-mini,qwq-32B; Standard models: deepseek-V3 (deepseekai2024deepseekv3technicalreport), llama-3.3-70B-Instruct, llama-3.2-3B-Instruct, qwen3-32B. The temperaturetemperature is $0.1$ where applicable (standard models), and the max_tokensmax\_tokens is $2046$.  [^7] Details on LLM configurations are in Appendix A.1.
Figure 3: F1 Performance as a function of training size. (A) Comparison of model architectures with BERT embeddings on Base structure organization. (B) Impact of data organization using the best architecture.
We test zero, one and five-shot prompting with (w-CoT) and without (wo-CoT) chain-of-thought reasoning (wei2021finetuned; kojima2022large). We use a puzzle-solving prompt that parallels our lightweight model setting. A zero-shot example of it is in Figure 2. The LLM first sees the context, generates a provisional answer, then compares it to the answer set before generating the best final sentence (instead of answer key label), and it also generates the hypotheses considered. Any final choice not in the answer set is flagged as a system error (ERR). This prompt mimics our learning objective using pretrained models and presents a more challenging task than standard multiple-choice questions. Randomizing the options in the answer set and asking to generate sentences instead of option labels help debias LLM’s token bias (e.g., A/B/C/D) in multiple-choice-format questions (zheng2024large). This uniform puzzle-solving format enables direct comparisons across model scales and training paradigms.
@@
F1 scores balance precision and recall across 7 answer choices. Responses not matching answer set options are flagged as system errors for LLMs. Multi-factor ANOVA tests Structure ×× Model ×× Shots ×× CoT interactions ($p<0.05$).
@@ 3.3 Results
Figure 4: Isolated contributions of organizational components to best model performance. (A)  Impact of analogical organisation, comparing Base against NoAnalogy and Shuffled. (B)  Impact of implicit soft annotations, comparing Base with NoSoftCue and Transposed variants.
@@ 3.3.1 Sample Efficiency Validation
Figure 3 confirms our core finding: lightweight models with Base organization achieve $F1=0.95$ with only $100$ examples, outperforming our best zero-shot reasoning model GPT-o3 ($F1=0.87$). Performance stabilizes at 100012001000-1200 examples across architectures. Panel B shows the impact of data organization, where Base structure outperforms Shuffled and Transposed arrangements. Identical content produces different learning trajectories, confirming that organizational structure, not information quantity, drives efficiency gains.
@@ 3.3.2 Component Analysis
Figure 4 validates systematic contributions from individual organizational components. While all structures eventually converge to similar performance levels at larger training sizes, analogical structure (Base vs. NoAnalogy vs.Shuffled) shows slightly more consistent advantages in the 5001,500500-1,500 example range where data efficiency matters most. Contextual cues (Base vs. NoSoftCue vs. Transposed) provide additional but smaller benefits.
This subtle differences in learning curves suggest that while both components contribute to learning, modern neural architectures can compensate for deficiencies in either component as training sizes increases.
@@ 3.3.3 LLM Comparison
Figure 5 confirms our third contribution: lightweight models trained on $1000$ structured examples outperform all tested LLMs in zero-shot conditions, including state-of-the-art reasoning models. Multi-factor ANOVA reveals significant main effects for Structure, Model, and Shots, with minimal CoT impact [^8] Complete results on Multi-factor ANOVA appear in Appendix D.2 (Table 3). . This confirms this task engages pattern recognition and rule application rather than multi-step reasoning. [^9] Complete results on LLMs performance are in the Appendix D.3 (Table 4).
@@
Reasoning models start with high zero-shot performance ($F1=0.8+$) but plateau quickly, while standard models begin lower (F1=0.10.5F1=0.1-0.5) but show steeper improvement curves, making them more sensitive to organizational structure.
Organizational Structure Patterns
Shuffled consistently performs worst across all models, confirming that random organization disrupts learning. Base and NoAnalogy emerge as most effective, often performing comparably. Notably, NoAnalogy sometimes outperforms Base in zero-shot settings for reasoning models, suggesting analogical structure becomes less crucial when models already possess strong reasoning capabilities. However, standard models benefit most from organizational effects that have greater room for improvement.

### v2

@@ Abstract
Large language models achieve strong performance on many tasks, but their training makes it hard to see which properties of the input support efficient linguistic rule learning. We ask how three cognitively-inspired principles of input design support sample-efficient linguistic rule induction: analogical structure, contrastive learning, and minimal contextual cue. We also ask how their effects compare to those of LLMs on the same controlled tasks. We implement these principles in structured sentence completion tasks that test English verb alternations. Lightweight models trained on hundreds to one-thousand such examples learn the alternation rules with high F1 on these tasks. Ablation studies show that analogical organisation is the main driver of sample efficiency, and contrastive distractors and minimal context help further gains. We also evaluate zero- and few-shot LLMs on the same tasks. In this controlled setting, the lightweight models reach higher F1 with far fewer task-specific data. We treat this contrast as a comparison between learning regimes rather than a general verdict on LLMs. Our results show that careful input organisation supports sample-efficient learning of linguistic rules and reveals distinct learning signatures for trained lightweight models and prompted LLMs.
@@ 1 Introduction
Analogical reasoning, recognising parallelism of structural relationships across different contexts, enables efficient learning by abstracting patterns from minimal examples and transferring knowledge to novel situations (Lake et al., 2017). However, this principle has faced persistent scalability challenges across computational approaches: early symbolic systems required extensive knowledge engineering (Forbus et al., 1995), neural word embeddings showed inconsistent performance (Mikolov et al., 2013), and even transformer-based models like BERT show mixed capabilities (Ushio et al., 2021; Thrush et al., 2020). While recent systems achieve strong performance (Yasunaga et al., 2024; Jiayang et al., 2023; Wijesiriwardene et al., 2023), they still require substantial computational resources or show inconsistent results across complexity levels.
Rather than scaling analogical reasoning systems, we operationalise analogical principles through strategic input organisation. Research shows that processing constraints can optimise attention allocation (Christiansen and Chater, 2016), while contrastive learning frameworks show how systematic positive-negative comparisons improve discriminative learning (Chen et al., 2020; He et al., 2020). This suggests that structural organisation, not architectural scaling, may unlock analogical efficiency.
@@
In this paper, we ask how input organisation can support sample-efficient linguistic rule induction. We also ask how the behaviour of lightweight models that use such organisation compares to that of large language models on the same controlled tasks. We develop a computational approach that organises input data into analogical paradigms through three cognitively-inspired strategies: (i) Analogical Structure, the systematic paradigmatic relationships that support pattern recognition; (ii) Contrastive Learning, the systematically designed distractors that enable discriminative learning; and (iii) Minimal Contextual Cues, the implicit annotations providing essential semantic information without explicit labelling.
We evaluate this setup using English verb alternations (Levin, 1993) through structured completion tasks based on Blackbird Language Matrices (Merlo, 2023). These tasks require identifying correct sentence completions from systematically organised contexts with contrastive alternatives. Figure 1 illustrates our approach: contexts encode analogical mappings across parallel paradigms (e.g., Man:Dice :: Exploring:Mat), requiring models to abstract structural and semantic role relationships rather than memorise surface forms.
Results demonstrate sample efficiency: lightweight models (BERT+CNN, $\sim 0.5M$ parameters) trained on only $100$ structured examples achieve $F1=0.95$, and, in this setting, is higher than zero-shot GPT-o3 ($F1=0.87$) on the same tasks. Cross-phenomenon validation confirms robustness beyond causative constructions.
This paper makes three contributions. First, we introduce a computational approach that organises input into analogical paradigms for sample-efficient linguistic rule induction. Second, we provide empirical evidence that this approach enables competitive performance for lightweight models with minimal training examples. Third, we show that, in this controlled setting, lightweight models reach higher F1 than zero-and few-shot LLMs on the same tasks, under a different learning regime.
In all experiments, lightweight models are fine-tuned on 1001000100-1000 task-specific examples. LLMs are evaluated in zero-shot and few-shot prompt-based settings without parameter updates. We therefore treat this as a comparison between learning regimes rather than a general verdict on LLM capabilities.
@@ 2 Data and Task Design
Our data consist of structured sentence completion tasks based on Blackbird Language Matrices  (Merlo, 2023). We extend this framework to instantiate three cognitively inspired principles of input design. This section describes our task framework, linguistic domain, and distractor design methodology.
@@ 2.1 Task Framework
Our completion tasks organise linguistic data into analogical paradigm structures. Each instance contains a structured Context comprising two parallel paradigms encoding analogical relationships (Pattern A $\leftrightarrow$ Pattern B), supplemented with minimal contextual cues that provide semantic scaffolding without explicit labelling. It also contains a systematic Answer Set: One correct completion that maintains established patterns, and six systematically designed distractors that violate specific structural, semantic, or paradigmatic constraints.
Success requires integrating our three cognitive principles. Models must recognise analogical mappings across paradigms (analogical structure), discriminate between valid and invalid transformations (contrastive learning), and use semantic scaffolding from minimal soft annotations (minimal contextual cues).
@@ 2.2 Linguistic Test Domain: English Verb Alternations
We validate the generalisability of the approach with bake-class verbs, a class of verbs exhibiting the unspecified object alternations (The chef baked a cake vs. The chef baked). In this class, the subject of both the transitive and the intransitive is an Agent. While the two classes share syntactic structure, they instantiate semantically distinct processes, testing whether our organisational principles generalise beyond specific semantic classes.
2.3 Analogical Paradigm Organisation
Figure 1 illustrates our systematic organisation approach. Each context follows a $2\times 4$ paradigmatic structure.
@@ 2.3 Analogical Paradigm Organisation
We provide semantic scaffolding through subtle annotations within each paradigm. For example, action descriptors indicate agentivity (The player did it) while state descriptions suggest thematic roles (The ball was on the floor). These cues clarify semantic relationships without explicit grammatical labelling.
@@
Each distractor violates exactly one constraint dimension while preserving others, enabling precise evaluation of rule component acquisition. Role errors test semantic understanding, structural errors test syntactic knowledge, and paradigm errors test analogical consistency across contexts.
@@
2.4 Data Generation
Our systematic generation process creates sentences based on templates with controlled lexical variation and seed alternation pairs, following established verb classifications (Levin, 1993). [^1] Verb class and seed sentences are detailed in Appendix A. We follow a hybrid, expert-guided process. Linguistic experts select and validate seed sentences and design a small set of templates with soft annotations for each alternation class. We then automatically instantiate full paradigms from these seeds and templates, controlling lexical variation and error types. This procedure creates a dataset of structured completion tasks over roll- and bake-class verbs, which we refer to as BLM-Roll-BakeE.  [^2] Data is available at https://www.idiap.ch/en/scientific-research/data/blm-roll-bakee/
@@ 3 Experiments
We systematically evaluate our analogical paradigm organisation approach through controlled experiments testing three core claims: (i) our organisational principles enable sample efficiency compared to unstructured baselines; (ii) individual components contribute systematically to performance; (iii) lightweight models trained with structured data can match zero-shot and few-shot LLMs on the same controlled linguistic rule learning tasks, under a different learning regime. The performance on BLM puzzle completion is used as a proxy for a model’s level of successful linguistic rule learning.
@@ 3.1 Experimental Design
We test five systematic conditions that isolate organisational components while controlling information content.
Base is the complete paradigmatic organisation with analogical structure and contextual cues. Shuffled has identical content in random order, (tests organisational vs. content effects). NoAnalogy removes the first paradigm (tests analogical structure contribution). NoSoftCue removes contextual annotations (tests scaffolding contribution). Transposed preserves analogical patterns but changes spatial arrangement ( $C_{\text{Trans}}=C_{\text{Base}}^{\top}$).  [^3] See the formal definitions in Appendix B.1.
@@
Type I uses the same verbs across paradigms ($3000$ examples, 80:10:10 splits, direct analogical mapping). Type II uses different verbs across paradigms ($15000$ examples, same splits, tests abstraction). Cross-phenomenon uses unspecified object alternations with bake-class verbs (validates generalisability).
@@
Training sizes range from $10$ to $2700$ examples for lightweight models; and $300$-example test sets (same as for Type I test data, and $300$ random examples from Type II test set) for LLMs. This setup balances computational feasibility with statistical reliability for cross-model comparisons.
@@ 3.2 Models and Training
We use several models of representation. We embed each sentence using BERT444bert-base-multilingual-cased(Devlin et al., 2019) as our primary model, RoBERTa555xlm-roberta-base (Liu et al., 2019) and ELECTRA666electra-base-discriminator (Clark et al., 2020) for encoder comparison. We reason that using a weaker base clarifies contributions attributable to data organisation rather than encoder capabilities.
As for architecture types, we use previously tested parameter-efficient neural models (CNN and FFNN in Nastase and Merlo (2023) for comparison). The CNN captures localised patterns and positional relationships, while the FFNN processes concatenated embeddings through fully-connected layers, integrating distributed information across the entire context. Both architectures maintain minimal parameter counts ($\sim 0.5M$ parameters) compared to the encoder ($\sim 110M$), isolating the influence of data organisation from architectural capacity. Both models output an answer embedding given the concatenated embeddings of the sentences from the context, evaluated against answer candidates using a max-margin objective with cosine similarity. Training runs for $120$ epochs, with learning rate $0.001$, Adam Optimizer, batch size $100$, and early stopping (patience $10$). All experiments use random seed $42$ with results averaged over $3$ independent runs. [^7] Full encoder comparison and replication details are provided in Appendix C.
@@
We evaluate eight advanced models including reasoning-capable systems.
Reasoning models: deepseek-R1 (DeepSeek-AI, 2025), gpt-o3, gpt-o3-mini,qwq-32B.
Standard models: deepseek-V3 (DeepSeek-AI, 2024), llama-3.3-70B-Instruct, llama-3.2-3B-Instruct, qwen3-32B.
The temperaturetemperature is $0.1$, where applicable (standard models), and the max_tokensmax\_tokens is $2046$.  [^8] Details on LLM configurations are in Appendix D.1.
Figure 3: F1 Performance as a function of training size. (A) Comparison of model architectures with BERT embeddings on Base structure organisation. (B) Impact of data organisation using the best architecture.
Prompting and Evaluation
We test zero, one and five-shot prompting with (w-CoT) and without (wo-CoT) chain-of-thought reasoning (Wei et al., 2021; Kojima et al., 2022). We use a puzzle-solving prompt that parallels our lightweight model setting. Figure 2 shows a zero-shot example. The LLM first sees the context, generates a provisional answer, then compares it to the answer set before generating the best final sentence (instead of answer key label), and it also generates the hypotheses considered. Any final choice not in the answer set is flagged as a system error (ERR). This prompt mimics our learning objective using pretrained models and presents a more challenging task than standard multiple-choice questions. Randomising the options in the answer set and asking to generate sentences instead of option labels help debias LLMs’ token bias (e.g., A/B/C/D) in multiple-choice-format questions (Zheng et al., 2024). This uniform puzzle-solving format enables direct comparisons across model scales and training paradigms.
@@
F1 scores balance precision and recall across 7 answer choices. Responses not matching the answer set options are flagged as system errors for LLMs. A multi-factor ANOVA tests Structure $\times$ Model $\times$ Shots $\times$ CoT interactions ($p<0.05$).
@@ 3.3 Results
Figure 4: Isolated contributions of organisational components to best model performance. (A)  Impact of analogical organisation, comparing Base against NoAnalogy and Shuffled. (B)  Impact of implicit soft annotations, comparing Base with NoSoftCue and Transposed variants.
@@ 3.3.1 Sample Efficiency Validation
Figure 3 confirms our core finding: lightweight models with Base organisation achieve $F1=0.95$ with only $100$ examples, and, in this setting, are higher than our best zero-shot reasoning model GPT-o3 ($F1=0.87$) on the same tasks. Performance stabilises at 100012001000-1200 examples across architectures. Panel B shows the impact of data organisation, where Base structure outperforms Shuffled and Transposed arrangements. Identical content produces different learning trajectories, confirming that organisational structure, not information quantity, drives efficiency gains.
@@ 3.3.2 Component Analysis
Figure 4 validates the systematic contributions from individual organisational components. While all structures eventually converge to similar performance levels at larger training sizes, analogical structure (Base vs. NoAnalogy vs.Shuffled) shows slightly more consistent advantages in the 5001500500-1500 example range where data efficiency matters most. Contextual cues (Base vs. NoSoftCue vs. Transposed) provide additional but smaller benefits.
These subtle differences in learning curves suggest that while both components contribute to learning, modern neural architectures can compensate for deficiencies in either component as training sizes increases.
@@ 3.3.3 LLM Comparison
Figure 5 confirms our third contribution: lightweight models trained on $1000$ structured examples reach higher $F1$ than all tested LLMs in zero-shot conditions, including state-of-the-art reasoning models, on these tasks and under our training regime. Multi-factor ANOVA reveals significant main effects for structure, model, and number of shots, with minimal CoT impact [^9] Complete results on Multi-factor ANOVA appear in Appendix D.2 (Table 3). . This result suggests that this task engages pattern recognition and rule application rather than multi-step reasoning. [^10] Complete results on LLMs performance are in Appendix D.3 (Table 4).
@@
Reasoning models start with high zero-shot performance ($F1=0.8+$) but plateau quickly, while standard models begin lower (F1=0.10.5F1=0.1-0.5) but show steeper improvement curves, making them more sensitive to organisational structure.
Organisational Structure Patterns
Shuffled consistently performs worst across all models, confirming that random organisation disrupts learning. Base and NoAnalogy emerge as most effective, often performing comparably. Notably, NoAnalogy sometimes outperforms Base in zero-shot settings for reasoning models, suggesting analogical structure becomes less crucial when models already possess strong reasoning capabilities. However, standard models benefit most from organisational effects that have greater room for improvement.